Maria has set a goal to save $ 10 000 in a savings account that earns 2.4% compounded annually. How much must she deposit each year for five years if the deposits are made at the end of each year?
Answer by mananth(16949)

use the formula for the future value of a series of payments, compounded annually.
The formula is:
FV = P * ((1 + r)^n - 1) / r
FV is the future value of the annuity =$10,000),
P is the annual deposit,
r is the annual interest rate =0.024
n is the number of years 5 years
Rearrange the formula for P
P = FV * r / ((1 + r)^n - 1)
Plug the values
P = 10000 * 0.024 / ((1 + 0.024)^5 - 1
Use calculator to find the annual deposit

)